-3 3 25 A= 5 7 and B= 4. 1 (a) Find AB if possible. (b) Find BA if possible. (c) Find A’ if possible. (Hint: A² = A : A.)
Answer #1

Las 74] 3 1 1-(.*)-( * AB 2 -3 3 57 AB= -6 +20 6+ 5 15+23 15+7 V AB= 13 22 +3+ LIĆ BA= -3 25 st -6 + 15 - 15+21 BA: 8 ts 20 +please provide feedback:)